OpenStudy (anonymous):

Rewrite The equation so that x is a function of y. Show Your work 1)5y+x=-3+4y 2)3x-2y=4+7x 3)4y-3(x-2)=4x+8

OpenStudy (anonymous):

1) 5y+x=-3+4y -5y from both sides gives x=-3-y

OpenStudy (anonymous):

2) 3x-2y=4+7x add 2y to both sides 3x=4+7x+2y subtract 7x -4x=4+2y divide by 2 -2x=2+y divide again but by -2 x=-1-y/2

OpenStudy (anonymous):

3) 4y-3(x-2)=4x+8 4y-3x+6=4x+8 (distributed) add 3x and subtract 8 from both sides 4y-2=7x divide by 7 (4y-2)/7=x